OpenStudy (anonymous):

Travis has a savings account that his parents opened for him. It pays 6% annual interest. His uncle also opened and account for him, but it pays 8% annual interest. If there is $800 more in the account that pays 6%, and the total annual interest from both accounts is $104, how much money is in each of the accounts? Show your work, using algebra.

OpenStudy (anonymous):

x is the money in 6% account. y is the money in 8% account. there's $800 more in the 6% account: x = y + 800 the total annual interest is $104: 0.06x + 0.08y = 104 Now substitute "y + 800" for x in the second equation and solve for y. Then add 800 and you have x.

OpenStudy (anonymous):

did you get it? or you need more explanation on this?

OpenStudy (anonymous):

I didn't the answer...do i subsitute y+800 and then after that what do I do

OpenStudy (anonymous):

you substitute (y+800) for x in the equation 0.06x + 0.08y = 104 in which it is when u substitute it looks like this 0.06(Y+800) + 0.08y = 104 and then solve for y.

OpenStudy (anonymous):

0.08 y+0.06 (y+800) = 104 Expand out terms of the left hand side: 0.14 y+48. = 104 Subtract 48. from both sides: 0.14 y = 56. Divide both sides by 0.14: y = 400.